1/3x + 1/2y + 1/2z = -1
1/5x - 1/3y + 1/4z = -1/3
1/2x + 1/ y + 1/2z = -1/2

Found 2 solutions by FrankM, Alan3354:
Answer by FrankM(1040)

The coefficient of y in last equation is missing. Please edit, and resubmit.

Answer by Alan3354(69443)

1/3x + 1/2y + 1/2z = -1
---------
Is it 1/(3x) ?
Or (1/3)x ?
===================
1/5x - 1/3y + 1/4z = -1/3
1/2x + 1/ y + 1/2z = -1/2